OpenStudy (anonymous):

Two asteroids are 75,000 m apart. One has a mass of 8 107 kg. If the force of gravity between them is 1.14 N, what is the mass of the other asteroid? A. 3.4 1011 kg B. 1.2 1012 kg C. 8.3 1012 kg D. 1.2 1010 kg

OpenStudy (abmon98):

Use Newtons law of Universal Gravitation Force of Gravity=gm1m2/(r)2 G is a constant and never changes and is equal to 6.67*10^-11 1.14=6.67*10^-11*8107*m2/(75000)^2 1.14*(75000)^2/6.67*10^-11*8107=m2